# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 4.1 Missing Data in Social Science
Data in social science often suffers from missing values (Kalton and Kasprzyk, 1982; Little, 1988; Little and Rubin, 1989). For instance, the answers in questionnaires may be incomplete due to the refusal of respondents to certain questions. Fig. 3 shows an example of questionnaire with missing values. Indeed, as people usually face numerous spam emails and pop-up advertisements and pay more attention to their privacy, it becomes increasingly difficult for social scientists to collect complete surveys. Little and Rubin (1989) reviewed the methods of handling missing data in social science datasets and presented three strategies. The first strategy is imputation that replaces missing values with suitable estimates and then performs standard complete-data methods on the filled-in data. The second strategy is weighting that drops the incomplete cases and assigns new weights to the complete cases to compensate for the information loss. The last strategy is direct analysis of the incomplete data that, for example, calculates the covariances using only observed pairs of variables. The review by Porter and Ecklund (2012) emphasized the importance of understanding the mechanisms driving missing data patterns.
It is worth mentioning that a questionnaire often contains multiple types of data, such as numerical data, categorical data, ordinal data, and even unstructured text. Missing data imputation for categorical data, ordinal data, and text are much more difficult than that for numerical data (Quintero and LeBoulluec, 2018). In Section 6.5, we will detail the imputation methods for non-numerical data.
@@ 4.2 Missing Data in Bioinformatics
Bioinformatics is an interdisciplinary field that applies methods from biology, computer science, and statistics to analyze biological data such as gene expression and network data. However, the prevalence of missing values in such datasets often prevents the direct application of standard analytical methods. For instance, DNA microarrays—a classical technology for gene expression profiling—frequently contain missing values due to various factors, including insufficient resolution or image corruption. The need for imputation in this context is well-documented. For instance, Troyanskaya et al. (2001) reviewed methods for DNA microarrays and demonstrated the effectiveness of k-nearest neighbor (kNN) and singular value decomposition approaches. Similarly, Moorthy et al. (2014) provided a review of several imputation algorithms for microarray gene expression data.
In the past decades, gene expression profiling technology has evolved a lot. For example, RNA sequencing (RNA-Seq), compared with microarray, allows for full sequencing of the whole transcriptome and has been widely used in the community of bioinformatics. Particularly, single-cell RNA sequencing (scRNA-Seq) provides the expression profiles of individual cells and can be used to uncover rare or new cell types within a cell population. As shown in Figure 4, scRNA-Seq data often contains a lot of missing values. Missing data imputation for scRNA-Seq has become a hot topic for a few years (Van Dijk et al., 2018; Huang et al., 2018; Li and Li, 2018; Lopez et al., 2018; Linderman et al., 2018; Chen and Zhou, 2018; Gong et al., 2018; Huang et al., 2018; Eraslan et al., 2019; Mongia et al., 2019; Arisdakessian et al., 2019; Xu et al., 2020b; Mongia et al., 2020; Xu et al., 2020c; Wang et al., 2021a; Rao et al., 2021; Dai et al., 2022). These methods can be organized into two categories: model-based methods (e.g. Van Dijk et al. (2018); Huang et al. (2018); Li and Li (2018)) and machine learning-based methods (e.g. (Mongia et al., 2019, 2020; Arisdakessian et al., 2019; Wang et al., 2021a; Xu et al., 2020c; Rao et al., 2021)). Zhang and Zhang (2018) conducted comprehensive experiments to compare many computational methods for imputing scRNA-Seq data. Dai et al. (2022) provided a platform for comparing different missing data imputation methods for scRNA-Seq.
One challenge in scRNA-Seq imputation is that it is difficult to distinguish those genes not expressed at all from those produced by dropout (Miao et al., 2019).
@@ 4.3 Missing Data in Healthcare
In healthcare, a critical data source is the Electronic Health Record (EHR), a systematized collection of a patient’s health information in digital format. EHRs typically encompass a wide range of data, including personal statistics, medical history, vital signs, medications, allergies, immunization status, laboratory test results, and radiology images. Originally used primarily for documenting patient information, EHRs are now leveraged with statistical and machine learning methods to improve healthcare outcomes (Shickel et al., 2017; Yadav et al., 2018). For instance, Miotto et al. (2016) used deep learning to predict patients’ future diseases from EHRs.
@@
Several studies have focused on missing data imputation for EHRs (Beaulieu-Jones et al., 2017, 2018; Xu et al., 2020a; Alamoodi et al., 2021; Kabir and Farrokhvar, 2022). For example, Beaulieu-Jones et al. (2017) proposed a deep autoencoder to recover missing values. Zhang et al. (2020b) used Extreme Gradient Boosting (XGBoost) (Chen et al., 2015) to predict missing values in the MIMIC-III dataset (Johnson et al., 2016), a well-known, freely accessible clinical database. Zhou and Saghapour (2021) developed a graphical tool called ImputEHR for implementing various imputation methods, from simple techniques to sophisticated gradient-boosted tree-based and neural network approaches. Furthermore, Liu et al. (2023a) provided a review of deep learning-based imputation techniques for healthcare data, and Luo (2022) evaluated state-of-the-art imputation methods specifically for clinical data from MIMIC-III.
@@ 4.4 Missing Data in Image Science and Computer Vision
Unlike tabular data, images possess a strong local structure, meaning pixels within a small region or patch often have similar values. Furthermore, images, particularly hyperspectral images, exhibit channel-wise correlations. Consequently, effective imputation of missing values in images must leverage both the local and channel structure. Numerous image inpainting methods have been developed, which we categorize as follows: exemplar-based methods (Criminisi et al., 2004), total-variation based methods (Chan et al., 2006; Benning et al., 2013), dictionary learning based methods (Shi et al., 2011; Li et al., 2014), matrix completion based methods (Wen et al., 2012; Hu et al., 2012; Fan and Chow, 2018), and deep learning based methods (Xie et al., 2012a; Fan and Cheng, 2018; Xiang et al., 2023). It is noteworthy that most deep learning-based methods require a set of training images, whereas the other categories can typically handle single images without prior model training. More detailed literature reviews can be found in (Guillemot and Le Meur, 2013; Shen et al., 2015; Elharrouss et al., 2020; Qin et al., 2021; Jam et al., 2021; Xiang et al., 2023).
Video inpainting Huang et al. (2016); Patwardhan et al. (2007); Yu et al. (2018); Xu et al. (2019b) involves replenishing missing segments in a video sequence with content that maintains spatial and temporal coherence. Also known as video completion, this technique has numerous practical applications, such as the removal of unwanted objects and video restoration. A recent survey by Quan et al. (2024) comprehensively reviews deep learning methods for image and video inpainting.
@@ 4.5 Missing Data in E-commerce and Social Media
Recommendation systems are generally categorized into three types (Breese et al., 1998; Sarwar et al., 2001; Adomavicius and Tuzhilin, 2005; Zhang et al., 2019a): content-based methods, collaborative filtering, and hybrid methods. Content-based methods recommend items that are similar to those a user has liked, or target users who are similar to those who have liked an item. This similarity is typically derived from side information, such as item attributes (e.g., genre) or user profiles (e.g., sex, age, and occupation). In contrast, collaborative filtering leverages latent correlations among users and items to predict unknown preferences. Hybrid methods, as the name suggests, combine content-based and collaborative filtering approaches. Notably, collaborative filtering is closely related to missing data imputation, while content-based methods more closely resemble regression with incomplete response variables. Given this paper’s focus, we will concentrate primarily on collaborative filtering (CF).
@@
We simply organize CF methods into the following four categories: similarity-based methods (Resnick et al., 1994; Fkih, 2022; Sarwar et al., 2002; Beregovskaya and Koroteev, 2021), matrix factorization-based methods, neural network-based methods, and other methods (Shani et al., 2002; Breese et al., 1998; Miyahara et al., 2000). Particularly, in the past decades, matrix factorization-based methods (Billsus and Pazzani, 1998; Mnih and Salakhutdinov, 2008; Koren et al., 2009; Fan et al., 2019; Rendle et al., 2019) have been extensively studied and used in CF. These methods usually exploit the potential low-rank structure of the incomplete user-item utility matrix via embedding items and users into a latent space of reduced dimension, where the observed utilities are approximated by the inner products of the user feature vectors and item feature vectors. More details will be introduced in Section 5.5.1. In recent years, neural networks and deep learning have shown promising performance in CF (Salakhutdinov et al., 2007; Dziugaite and Roy, 2015; Sedhain et al., 2015; Wu et al., 2016; Zheng et al., 2016; He et al., 2017; van den Berg et al., 2017; Muller et al., 2018; Fan and Cheng, 2018; Yi et al., 2020). For instance, Sedhain et al. (2015) proposed an autoencoder (Hinton and Salakhutdinov, 2006) based CF method called AutoRec. It predicts unknown ratings by an encoder-decoder model, though the input of the encoder contains missing values (usually filled with zeros). AutoRec outperformed a few matrix factorization-based methods, such as SVD++ (Koren et al., 2009) and LLORMA (Lee et al., 2016), on several benchmark datasets (Sedhain et al., 2015). More about collaborative filtering can be found in the following review papers (Burke, 2002; Su and Khoshgoftaar, 2009; Bobadilla et al., 2013; Wang et al., 2021c).
@@ 4.6 Missing Data in Manufacturing Industries
In manufacturing industries, such as chemicals and consumer electronics, failures in numerous equipment and sensors can lead to missing data (Lakshminarayan et al., 1999; Imtiaz and Shah, 2008; Ehrlinger et al., 2018). The data in these contexts are often multivariate time series, making it essential to account for temporal dynamics when recovering missing values. Imtiaz and Shah (2008) focused on process data and categorized missing data patterns in chemical processes into four types (shown in Figure 7): (a) sensor breakdown; (b) process shutdown; (c) general patterns (e.g., outliers combined with sensor breakdown); and (d) multi-rate data. These patterns are typically classified as Missing Not at Random (MNAR).
Several studies have addressed missing data problems in manufacturing industries (Imtiaz and Shah, 2008; Ziaei-Halimejani et al., 2021; Xie et al., 2020; Carbery et al., 2022; Fan et al., 2022; Jeong et al., 2023). For example, Ehrlinger et al. (2018) described the challenges associated with missing data in industrial data analytics, while Du et al. (2020) reviewed the missing data problem in sensor network-based monitoring systems.
@@ 5.1 Basic Imputation via Filling with Zero, Mean, Median, or Mode
Filling missing values with zeros often yields statistical bias because the means of the variables are not necessarily zero. Note that in count data such as scRNA-Seq data, a zero could be either an observed value or a missing value, which means in many scenarios the missing values are naturally filled with zeros. On the other hand, zero filling can be a useful initialization for some missing data imputation algorithms such as low-rank matrix completion (see Section 5.5.1) and autoencoder based imputation (see Section 5.6). An appealing property of zero filling is that for regression models, especially neural networks, the contribution of the corresponding missing value is prohibited, which may improve the robustness or generalization ability of the models (Fan et al., 2024). A few researcher showed that zero filling has an adverse effect on model performances (Śmieja et al., 2018; Yi et al., 2019).
@@ 5.2 Regression Imputation
The aforementioned zero, mean, median, and mode imputations only consider the distribution of the values of the variable that has missing entries. However, if there is a correlation between the target variable and other variables, which is often the case in practice, regression of the missing variable on these other variables often provides more accurate estimations especially when data are MNAR (Little, 1988; Roth, 1994). For instance, in a bivariate analysis where one variable has missing data, the regression equation is estimated using only the cases with complete data. In this equation, the variable with incomplete data is treated as the outcome, while the variable with complete data serves as the predictor. Baraldi and Enders (2010) provided an intuitive example that used math score to predict course grade. More generally, suppose we have $d$ variables $x_{1},x_{2},\ldots,x_{d}$ and we want to impute the missing values for $x_{j}$. Regression imputation first uses the complete observations to build a regression model, such as
@@
where $\varepsilon$ denotes the error term and $f^{(j)}$ could be a linear function or nonlinear function. An example of $f^{(j)}$ is $\sum_{i\neq j}w_{i}x_{i}$, where $w_{i}$ are the coefficients.
An important variation of regression imputation is the iterative regression method (Roth, 1994). It begins by calculating an initial correlation matrix using a straightforward imputation method, like mean imputation, or simply using the observed values. Following this, regression equations are derived from this matrix, and the missing values are estimated and replaced in the data matrix. This cycle continues until the regression coefficients exhibit minimal changes.
Perhaps the best-known iterative regression imputation algorithm is the missForest proposed by Stekhoven and Bühlmann (2011). In missForest, the $f^{(j)}$ in (1) is a random forest, and there are $d$ random forests in total. The main steps of missForest are summarized as follows:
@@
One of the advantages of missForest is that it can handle non-linear relationships and interactions between variables, since $f^{(j)}$ are random forests. It also provides an out-of-bag estimate of the imputation error, which can be useful for assessing the quality of the imputations. In addition, missForest is effective in handling mixed-type data, including numerical and categorical variables.
The random forest within the missForest algorithm can be replaced with other regression models, such as polynomial regression, support vector regression (Drucker et al., 1996), Gaussian processes (Williams and Rasmussen, 1995), or neural networks (LeCun et al., 2015). This substitution can create new regression-based imputation algorithms, though potentially at the cost of higher computational complexity and reduced effectiveness with categorical data. In practice, missForest remains a stable and effective benchmark. To the best of our knowledge, no variant proposed in the past decade has substantially surpassed its imputation accuracy. For instance, while Liu and Constantinou (2023) incorporated a Markov Blanket discovery method to refine the feature selection process in missForest, it achieved only a marginal improvement in accuracy.
Regression imputation has several potential drawbacks. It can artificially reduce the variance of imputed variables and introduce bias if the underlying model assumptions are violated (Roth, 1994; Baraldi and Enders, 2010). Furthermore, if the proportion of missing data is very high (e.g., 90%), estimating the regression function $f^{(j)}$ may be infeasible or highly inaccurate, leading to imputation failure. Although iterative regression methods can mitigate this issue, they remain sensitive to the initial estimates, and an unreliable starting correlation matrix can still yield poor final results.

### v2

@@ 4.1 Missing Data in Social Science
Data in social science often suffers from missing values (kalton1982imputing; little1988missing; little1989analysis). For instance, questionnaire responses may be incomplete because respondents refuse to answer certain questions. Fig. 3 shows an example of a questionnaire with missing values. Indeed, as people usually face numerous spam emails and pop-up advertisements and pay more attention to their privacy, it becomes increasingly difficult for social scientists to collect complete surveys. little1989analysis reviewed methods for handling missing data in social science datasets, outlining three primary strategies. The first is imputation, which replaces missing values with appropriate estimates, allowing standard complete-data methods to be applied. The second is weighting, which involves dropping incomplete cases and assigning new weights to the complete cases to compensate for the lost information. The last strategy is direct analysis of the incomplete data that, for example, calculates the covariances using only observed pairs of variables. The review by porter2012missing emphasized the importance of understanding the mechanisms driving missing data patterns.
It is worth mentioning that a questionnaire often contains multiple types of data, such as numerical data, categorical data, ordinal data, and even unstructured text. Missing data imputation for categorical data, ordinal data, and text are much more difficult than that for numerical data (quintero2018missing). In Section 6.5, we will detail the imputation methods for non-numerical data.
@@ 4.2 Missing Data in Bioinformatics
Bioinformatics is an interdisciplinary field that applies methods from biology, computer science, and statistics to analyze biological data such as gene expression and network data. However, the prevalence of missing values in such datasets often prevents the direct application of standard analytical methods. For instance, DNA microarrays—a classical technology for gene expression profiling—frequently contain missing values due to various factors, including insufficient resolution or image corruption. The need for imputation in this context is well-documented. For instance, troyanskaya2001missing reviewed methods for DNA microarrays and demonstrated the effectiveness of k-nearest neighbor (kNN) and singular value decomposition approaches. Similarly, moorthy2014review provided a review of several imputation algorithms for microarray gene expression data.
In the past decades, gene expression profiling technology has evolved a lot. For example, RNA sequencing (RNA-Seq), compared with microarray, allows for full sequencing of the whole transcriptome and has been widely used in the community of bioinformatics. Particularly, single-cell RNA sequencing (scRNA-Seq) provides the expression profiles of individual cells and can be used to uncover rare or new cell types within a cell population. As shown in Figure 4, scRNA-Seq data often contains a lot of missing values. Missing data imputation for scRNA-Seq has become a hot topic for a few years (van2018recovering; huang2018saver; li2018accurate; lopez2018deep; linderman2018zero; chen2018viper; gong2018drimpute; huang2018saver; eraslan2019single; mongia2019mcimpute; arisdakessian2019deepimpute; xu2020cmf; mongia2020deepmc; xu2020scigans; wang2021scgnn; rao2021imputing; dai2022scimc). These methods can be organized into two categories: model-based methods (e.g. van2018recovering; huang2018saver; li2018accurate) and machine learning-based methods (e.g. (mongia2019mcimpute; mongia2020deepmc; arisdakessian2019deepimpute; wang2021scgnn; xu2020scigans; rao2021imputing)). zhang2018comparison conducted comprehensive experiments to compare many computational methods for imputing scRNA-Seq data. dai2022scimc provided a platform for comparing different missing data imputation methods for scRNA-Seq.
One challenge in scRNA-Seq imputation is that it is difficult to distinguish those genes not expressed at all from those produced by dropout (miao2019screcover).
@@ 4.3 Missing Data in Healthcare
In healthcare, a critical data source is the Electronic Health Record (EHR), a systematized collection of a patient’s health information in digital format. EHRs typically encompass a wide range of data, including personal statistics, medical history, vital signs, medications, allergies, immunization status, laboratory test results, and radiology images. Originally used primarily for documenting patient information, EHRs are now leveraged with statistical and machine learning methods to improve healthcare outcomes (shickel2017deep; yadav2018mining). For instance, miotto2016deep used deep learning to predict patients’ future diseases from EHRs.
@@
Several studies have focused on missing data imputation for EHRs (beaulieu2017missing; beaulieu2018characterizing; XU2020103576; ALAMOODI2021111236; kabir2022non). For example, beaulieu2017missing proposed a deep autoencoder to recover missing values. zhang2020predicting used Extreme Gradient Boosting (XGBoost) (chen2015xgboost) to predict missing values in the MIMIC-III dataset (johnson2016mimic), a well-known, freely accessible clinical database. zhou2021imputehr developed a graphical tool called ImputEHR for implementing various imputation methods, from simple techniques to sophisticated gradient-boosted tree-based and neural network approaches. Furthermore, liu2022handling provided a review of deep learning-based imputation techniques for healthcare data, and luo2022evaluating evaluated state-of-the-art imputation methods specifically for clinical data from MIMIC-III.
@@ 4.4 Missing Data in Image Science and Computer Vision
Unlike tabular data, images possess a strong local structure, meaning pixels within a small region or patch often have similar values. Furthermore, images, particularly hyperspectral images, exhibit channel-wise correlations. Consequently, effective imputation of missing values in images must leverage both the local and channel structure. Numerous image inpainting methods have been developed, which we categorize as follows: exemplar-based methods (1323101), total-variation based methods (chan2006total; benning2013higher), dictionary learning based methods (5995592; LI20142862), matrix completion based methods (wen2012solving; hu2012fast; FANNLMC), and deep learning based methods (NIPS2012_6cdd60ea; fan2018matrix; XIANG2023109046). It is noteworthy that most deep learning-based methods require a set of training images, whereas the other categories can typically handle single images without prior model training. More detailed literature reviews can be found in (guillemot2013image; shen2015missing; elharrouss2020image; qin2021image; jam2021comprehensive; XIANG2023109046).
Video inpainting 10.1145/2980179.2982398; patwardhan2007video; yu2018generative; xu2019deep involves replenishing missing segments in a video sequence with content that maintains spatial and temporal coherence. Also known as video completion, this technique has numerous practical applications, such as the removal of unwanted objects and video restoration. A recent survey by quan2024deep comprehensively reviews deep learning methods for image and video inpainting.
@@ 4.5 Missing Data in E-commerce and Social Media
Recommendation systems are generally categorized into three types (breeseCF1998; sarwar2001item; 2005CFsurvey; Zhang2019): content-based methods, collaborative filtering, and hybrid methods. Content-based methods recommend items that are similar to those a user has liked, or target users who are similar to those who have liked an item. This similarity is typically derived from side information, such as item attributes (e.g., genre) or user profiles (e.g., sex, age, and occupation). In contrast, collaborative filtering leverages latent correlations among users and items to predict unknown preferences. Hybrid methods, as the name suggests, combine content-based and collaborative filtering approaches. Notably, collaborative filtering is closely related to missing data imputation, while content-based methods more closely resemble regression with incomplete response variables. Given this paper’s focus, we will concentrate primarily on collaborative filtering (CF).
@@
We simply organize CF methods into the following four categories: similarity-based methods (knncf; FKIH20227645; sarwar2002recommender; beregovskaya2021review), matrix factorization-based methods, neural network-based methods, and other methods (markov1; breeseCF1998; bayes2). Particularly, in the past decades, matrix factorization-based methods (billsus98; Salakhutdinov07; koren09; fan2019factor; rendle2019difficulty) have been extensively studied and used in CF. These methods usually exploit the potential low-rank structure of the incomplete user-item utility matrix via embedding items and users into a latent space of reduced dimension, where the observed utilities are approximated by the inner products of the user feature vectors and item feature vectors. More details will be introduced in Section 5.5.1. In recent years, neural networks and deep learning have shown promising performance in CF (RBMCF; nnmf; Sedhain2015; CDAE; Zheng2016; he2017neural; Berg2017a; muller18a; fan2018matrix; Yi2020). For instance, Sedhain2015 proposed an autoencoder (hinton2006reducing) based CF method called AutoRec. It predicts unknown ratings by an encoder-decoder model, though the input of the encoder contains missing values (usually filled with zeros). AutoRec outperformed a few matrix factorization-based methods, such as SVD++ (koren09) and LLORMA (LLORMA), on several benchmark datasets (Sedhain2015). More about collaborative filtering can be found in the following review papers (burke2002hybrid; su2009survey; bobadilla2013recommender; wang2021survey).
@@ 4.6 Missing Data in Manufacturing Industries
In manufacturing industries, such as chemicals and consumer electronics, failures in numerous equipment and sensors can lead to missing data (lakshminarayan1999imputation; imtiaz2008treatment; 8846984). The data in these contexts are often multivariate time series, making it essential to account for temporal dynamics when recovering missing values. imtiaz2008treatment focused on process data and categorized missing data patterns in chemical processes into four types (shown in Figure 7): (a) sensor breakdown; (b) process shutdown; (c) general patterns (e.g., outliers combined with sensor breakdown); and (d) multi-rate data. These patterns are typically classified as Missing Not at Random (MNAR).
Several studies have addressed missing data problems in manufacturing industries (imtiaz2008treatment; ZIAEIHALIMEJANI2021107549; 8891716; carbery2022missingness; fantii2022; JEONG2023103937). For example, 8846984 described the challenges associated with missing data in industrial data analytics, while du2020missing reviewed the missing data problem in sensor network-based monitoring systems.
@@ 5.1 Basic Imputation via Filling with Zero, Mean, Median, or Mode
Filling missing values with zeros often yields statistical bias because the means of the variables are not necessarily zero. Note that in count data such as scRNA-Seq data, a zero could be either an observed value or a missing value, which means in many scenarios the missing values are naturally filled with zeros. On the other hand, zero filling can be a useful initialization for some missing data imputation algorithms such as low-rank matrix completion (see Section 5.5.1) and autoencoder-based imputation (see Section 5.6). An appealing property of zero filling is that for regression models, especially neural networks, the contribution of the corresponding missing value is prohibited, which may improve the robustness or generalization ability of the models (fan2024neuronenhanced). A few researchers showed that zero filling has an adverse effect on model performances (smieja2018processing; yi2019not).
@@ 5.2 Regression Imputation
The aforementioned zero, mean, median, and mode imputations only consider the distribution of the values of the variable that has missing entries. However, if there is a correlation between the target variable and other variables, which is often the case in practice, regression of the missing variable on these other variables often provides more accurate estimations, especially when data are MNAR (little1988missing; roth1994missing). For instance, in a bivariate analysis where one variable has missing data, the regression equation is estimated using only the cases with complete data. In this equation, the variable with incomplete data is treated as the outcome, while the variable with complete data serves as the predictor. baraldi2010introduction provided an intuitive example that used the math score to predict course grade. More generally, suppose we have $d$ variables $x_{1},x_{2},\ldots,x_{d}$ and we want to impute the missing values for $x_{j}$. Regression imputation first uses the complete observations to build a regression model, such as
@@
where $\varepsilon$ denotes the error term and $f^{(j)}$ could be a linear function or nonlinear function. An example of $f^{(j)}$ is $\sum_{i\neq j}w_{i}x_{i}+b$, where $w_{i}$ are the coefficients and $b$ is a bias term.
An important variation of regression imputation is the iterative regression method (roth1994missing). It begins by calculating an initial correlation matrix using a straightforward imputation method, like mean imputation, or simply using the observed values. Following this, regression equations are derived from this matrix, and the missing values are estimated and replaced in the data matrix. This cycle continues until the regression coefficients exhibit minimal changes.
Perhaps the best-known iterative regression imputation algorithm is the missForest proposed by MissForest. In missForest, the $f^{(j)}$ in (1) is a random forest, and there are $d$ random forests in total. The main steps of missForest are summarized as follows:
@@
One of the advantages of missForest is that it can handle non-linear relationships and interactions between variables, since $f^{(j)}$ are random forests. It also provides an out-of-bag estimate of the imputation error, which can be useful for assessing the quality of the imputations. Additionally, missForest is effective in handling mixed-type data, including numerical and categorical variables.
The random forest within the missForest algorithm can be replaced with other regression models, such as polynomial regression, support vector regression (drucker1996support), Gaussian processes (williams1995gaussian), or neural networks (lecun2015deep). This substitution can create new regression-based imputation algorithms, though potentially at the cost of higher computational complexity and reduced effectiveness with categorical data. In practice, missForest remains a stable and effective benchmark. To the best of our knowledge, no variant proposed in the past decade has substantially surpassed its imputation accuracy. For instance, while liu2023improving incorporated a Markov Blanket discovery method to refine the feature selection process in missForest, it achieved only a marginal improvement in accuracy.
Regression imputation has several potential drawbacks. It can artificially reduce the variance of imputed variables and introduce bias if the underlying model assumptions are violated (roth1994missing; baraldi2010introduction). Furthermore, if the proportion of missing data is very high (e.g., 90%), estimating the regression function $f^{(j)}$ may be infeasible or highly inaccurate, leading to imputation failure. Although iterative regression methods can mitigate this issue, they remain sensitive to the initial estimates, and an unreliable starting correlation matrix can still yield poor final results.